OpenStudy (anonymous):

5(x + 4) - 4x = 18 - 23

pooja195 (pooja195):

5(x+4)−4x=18−23 (5)(x)+(5)(4)+−4x=18+−23(Distribute) 5x+20+−4x=18+−23 (5x+−4x)+(20)=(18+−23) (Combine Like Terms) 5-4=?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

x=1 ?

pooja195 (pooja195):

well 5-4=1 20 remains the same we simplified the other side now we have x+20=−5 subtract 20 from BOTH sides to get your FINAL answer

OpenStudy (anonymous):

-25

OpenStudy (anonymous):

thx yet again obi wan